# Leveraging the Power of MLLMs for Gloss-Free Sign Language Translation

### v1

@@ 4.3 Sign Language Translation
To perform end-to-end gloss-free SLT, we first inherit the pre-trained networks from the MMLP. In this context, the description encoder $\psi_{de}$ is not utilized, and the description mapper $\psi_{dm}$ remains frozen, as it is solely employed for predicting SL descriptions. Hence, we fine-tune the visual encoder $\psi_{vis}$, the modality adapter $\psi_{ma}$, and the multimodal encoder $\psi_{enc}$. Given a sign language video ${SV}_{i}$, we extract the multimodal feature $\tilde{M}_{i}$. Subsequently, the initialized LLM decoder $\psi_{dec}$, which is structured as a 12-layer mBART decoder [48], takes the multimodal feature $\tilde{M}_{i}$ as input to produce the predicted spoken sentence $\hat{SL}_{i}=(\hat{SL}_{i,1},\dots,\hat{SL}_{i,\bar{T}})$.
@@ 5 Experiments
[…]
@@ 5.1 Datasets, Metrics, and Implementation Details
Datasets. Following prior studies [67, 12, 57, 18], we conduct experiments on two SLT benchmark datasets, PHOENIX14T [5] and CSL-Daily [68], evaluating our method on their development (dev) and test sets. PHOENIX14T is a German sign language dataset comprising three years of daily news and weather forecasts, divided into train, dev, and test sets containing 7,096, 519, and 642 videos, respectively. The German spoken sentences have a vocabulary size of 2,887. CSL-Daily is a large-scale continuous SLT dataset for Chinese Sign Language, covering various daily scenarios. The train, dev, and test sets contain 18,401, 1,077, and 1,176 videos, respectively, with a vocabulary size of 2,343 for Chinese spoken sentences.
@@
For implementation details, please refer to supplementary.
@@ 5.2 Comparison with State-of-the-Art
Results on CSL-Daily. Tab. 2 presents the results of the large-scale SLT benchmark dataset, CSL-Daily. Our model achieves SOTA performance across all metrics, demonstrating significant improvements of +5.36 in BLEU-4 and +6.56 in ROUGE scores over SignLLM [18] and Sign2GPT [57] on the test set, respectively. Surprisingly, it outperforms VAP [27], the SOTA weakly supervised gloss-free model, across all metrics. This improvement may be attributed to MLLM-generated SL descriptions, which capture SL component details more accurately in the high-resolution CSL-Daily dataset. A comparison of SL descriptions between PHOENIX14T and CSL-Daily is provided in the supplementary. These findings highlight the effectiveness of the synergy between MLLM-generated SL descriptions and sign language images in improving translation performance.
@@ 5.3 Ablation studies
In this section, to perform ablation studies, we conduct all experiments on the test set of PHOENIX14T.
Impact of Key Components in MMSLT. We conduct experiments to assess the influence of the key components constituting MMSLT, including the GSD-MLLM module, the multimodal-language alignment, and the description mapper. Tab. 3 presents the performance achieved by incrementally adding each component.
@@
A comparison between the first row, which translates using only sign language images, and the second row, which uses both sign language images and SL descriptions, indicates that the second row exhibits lower performance. This can be attributed to the modality gap between the sign language images and the corresponding SL descriptions. However, performance is substantially improved in the fourth row, where this gap is addressed through the multimodal-language alignment. Moreover, comparing the second and third rows demonstrates a significant performance gain when employing a description mapper that mitigates the input modality gap, even without alignment. This highlights the importance of effectively utilizing modules capable of addressing the modality gap when incorporating SL descriptions. Finally, a comparison between the fourth and final rows shows improved performance with the addition of the description mapper. This suggests that when combined with other modules, the description mapper creates a synergistic effect. These findings underscore that each component of MMSLT plays a significant role in contributing to performance improvement.
Quantitative Evaluation of MLLMs and Prompts. We evaluate MLLM-generated SL descriptions. Details on each MLLM’s characteristics are provided in the supplementary. First, to assess how effectively SL descriptions represent sign language images, we analyze the performance of using only SL descriptions without videos. Consistent with Sec. 3, the LLaVA-OneVision 7B model performs best, focusing on SL components. We also identify a limitation in using descriptions alone, as indicated by the deficient BLEU-4 score. Next, we analyze the results obtained using SL descriptions from various MLLMs within the proposed framework. All models except InternVL2 [11] outperform the previous gloss-free SOTA, SignLLM [18], highlighting consistent SLT improvements with SL descriptions.
@@
Prompts BLEU-1 BLEU-2 BLEU-3 BLEU-4
@@
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4
@@
Table 7: Visualization of translation results. Correct answers are marked in blue, answers with the same meaning but different wording are displayed in green, and incorrect answers are indicated in red.
5.4 Qualitative Results
Tab. 7 presents the translation results for randomly selected sign language videos from the PHOENIX14T test set. We visualize the reference, GFSLT-VLP [67] (the only other gloss-free model with publicly available code), and MMSLT. The results indicate that while GFSLT-VLP correctly translates only a few words or generates entirely different sentences, MMSLT performs comparatively competently, often replacing certain words with others of similar meaning while preserving the overall meaning. These examples qualitatively show MMSLT’s translation effectiveness.
@@ 6 Conclusion
In this study, we propose MMSLT, a gloss-free SLT framework that leverages an off-the-shelf MLLM for the first time. We utilize the MLLM to generate SL descriptions through carefully crafted prompts, integrating them with sign language images to effectively represent sign language. We introduce a pre-training module to align the fused modalities with spoken sentences, addressing modality gaps in SLT. Additionally, we propose a description mapper to reduce the computational burden of the MLLM during inference by approximating the SL descriptions. This research lays the foundation for using MLLMs in SLT and opens up new avenues for future exploration in the field.

### v2

@@ 4.3 Sign Language Translation
To perform end-to-end gloss-free SLT, we first inherit the pre-trained networks from the MMLP. In this context, the description encoder $\psi_{de}$ is not utilized, and the description mapper $\psi_{dm}$ remains frozen, as it is solely employed for predicting SL descriptions. Hence, we fine-tune the visual encoder $\psi_{vis}$, the modality adapter $\psi_{ma}$, and the multimodal encoder $\psi_{enc}$. Given a SL video ${SV}_{i}$, we extract the multimodal feature $\tilde{M}_{i}$. Subsequently, the initialized LLM decoder $\psi_{dec}$, which is structured as a 12-layer mBART decoder [48], takes the multimodal feature $\tilde{M}_{i}$ as input to produce the predicted spoken sentence $\hat{SL}_{i}=(\hat{SL}_{i,1},\dots,\hat{SL}_{i,\bar{T}})$.
@@ 5 Experiments
Table 2: Experimental results on CSL-Daily dataset. In VAP [27](marked with $\ddagger$), for a fair comparison, we present the performance without post-processing to convert punctuation to the Chinese format  [39] (e.g.,。,?) using Python’s replace function. $\ast$}, $\star$} represent the results reproduced by [62] and [67], respectively.
@@ 5.1 Datasets, Metrics, and Implementation Details
Datasets. Following prior studies [67, 12, 57, 18], we conduct experiments on two SLT benchmark datasets, PHOENIX14T [5] and CSL-Daily [68] evaluating our method on their development (dev) and test sets. PHOENIX14T is a German SL dataset comprising three years of daily news and weather forecasts, divided into train, dev, and test sets containing 7,096, 519, and 642 videos, respectively. The German spoken sentences have a vocabulary size of 2,887. CSL-Daily is a large-scale continuous SLT dataset for Chinese SL, covering various daily scenarios. The train, dev, and test sets contain 18,401, 1,077, and 1,176 videos, respectively, with a vocabulary size of 2,343 for Chinese spoken sentences.
@@
Implementation Details. Please refer to supplementary.
@@ 5.2 Comparison with State-of-the-Art
Results on CSL-Daily. Tab. 2 presents the results of the large-scale SLT benchmark dataset, CSL-Daily. Our model achieves SOTA performance across all metrics, demonstrating significant improvements of +5.36 in BLEU-4 and +6.56 in ROUGE scores over SignLLM [18] and Sign2GPT [57] on the test set, respectively. Surprisingly, it outperforms VAP [27], the SOTA weakly supervised gloss-free model, across all metrics. This improvement may be attributed to MLLM-generated SL descriptions, which capture SL component details more accurately in the high-resolution CSL-Daily dataset. A comparison of SL descriptions between PHOENIX14T and CSL-Daily is provided in the supplementary. These findings highlight the effectiveness of the synergy between MLLM-generated SL descriptions and SL images in improving translation performance.
@@ 5.3 Ablation studies
In this section, we conduct all ablation experiments using the test set of PHOENIX14T. For further ablation studies, please refer to the supplementary materials.
Impact of Key Components in MMSLT. We conduct experiments to assess the influence of the key components constituting MMSLT, including the GSD-MLLM module, the multimodal-language alignment, and the description mapper. Tab. 3 presents the performance achieved by incrementally adding each component. A comparison between the first row, which translates using only SL images, and the second row, which uses both SL images and SL descriptions, indicates that the second row exhibits lower performance. This can be attributed to the modality gap between the SL images and the corresponding SL descriptions. However, performance is substantially improved in the fourth row, where this gap is addressed through the multimodal-language alignment. In addition, a comparison of the third and fourth rows highlights the effectiveness of GSD-MLLM when paired with M-L Align, as it surpasses M-L Align alone. Moreover, comparing the second and fifth rows demonstrates a significant performance gain when employing a description mapper that mitigates the input modality gap, even without alignment. This highlights the importance of effectively utilizing modules capable of addressing the modality gap when incorporating SL descriptions. Finally, a comparison between the fourth and final rows shows improved performance with the addition of the description mapper. This suggests that when combined with other modules, the description mapper creates a synergistic effect. These findings underscore that each component of MMSLT plays a significant role in contributing to performance improvement.
@@
Quantitative Evaluation of MLLMs and Prompts. We evaluate MLLM-generated SL descriptions. Details on each MLLM’s characteristics are provided in the supplementary. First, to assess how effectively SL descriptions represent SL images, we analyze the performance of using only SL descriptions without videos, as shown in Tab. 4. Consistent with Sec. 3, the LLaVA-OneVision 7B model performs best, focusing on SL components. We also identify a limitation in using descriptions alone, as indicated by the deficient BLEU-4 score. Next, we analyze the results obtained using SL descriptions from various MLLMs within the proposed framework. All models except InternVL2 [11] outperform the previous gloss-free SOTA, SignLLM [18], highlighting consistent SLT improvements with SL descriptions.
@@
Prompts B-1 B-2 B-3 B-4
@@
Model B-1 B-2 B-3 B-4
@@
5.4 Qualitative Results
Tab. 7 presents the translation results for randomly selected SL videos from the PHOENIX14T test set. We visualize the reference, GFSLT-VLP [67] (the only other gloss-free model with publicly available code), and MMSLT. The results indicate that while GFSLT-VLP correctly translates only a few words or generates entirely different sentences, MMSLT performs more competently, often replacing certain words with others of similar meaning while preserving the overall meaning. These examples qualitatively show MMSLT’s translation effectiveness.
@@ 5.4 Qualitative Results
Table 7: Translation results. Correct answers are marked in blue, semantically similar but rephrased answers are displayed in green, and incorrect answers are indicated in red.
@@ 6 Conclusion
In this study, we propose MMSLT, a gloss-free SLT framework that, for the first time, leverages an off-the-shelf MLLM. We utilize the MLLM to generate SL descriptions through carefully crafted prompts, integrating them with SL images to effectively represent SL. We introduce a pre-training module to align the fused modalities with spoken sentences, addressing modality gaps in SLT. Additionally, we propose a description mapper to reduce the computational burden of the MLLM during inference by approximating the SL descriptions. This research not only lays the foundation for using MLLMs in SLT but also paves the way for future advancements in multimodal learning across various computer vision fields.
7 Acknowledgement
This work was supported by the National Research Foundation (NRF) grants funded by Korean Government (MSIT) (No. 2023R1A2C200337911 and No. RS-2023-00220762) and the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2025-02303870, Software Technology for Efficient Multimodal Visual Information Processing in High-Speed Spatial Interactions).